# A Wasserstein perspective of Vanilla GANs

### v1

@@ 7.3 Proofs for Section 4
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\mathsf{W}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})+2\inf_{W^{\prime}\in\mathcal{W}}\sup_{W\in\operatorname{Lip}(1,B)}\|W-W^{\prime}\|_{\infty}$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\mathsf{W}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})+2\inf_{W^{\prime}\in\mathcal{W}}\sup_{W\in\operatorname{Lip}(1,B)}\|W-W^{\prime}\|_{\infty}$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\mathsf{W}_{\mathcal{H}^{\alpha}(\Gamma)}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})+2\inf_{W^{\prime}\in\mathcal{W}}\sup_{W\in\operatorname{Lip}(1,B)}\|W-W^{\prime}\|_{\infty}.$
The bound on $\mathsf{W}_{\mathcal{H}^{\alpha}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})$ depending on the intrinsic dimension $d^{*}$ was already derived in Theorem 5.2. Along Corollary 4.2 we obtain the rate depending on the ambient dimension $d.$
Note that the Wasserstein distance is invariant to additive shifts, hence we can add the assumption $W(0)=0$ without changing the optimal value. As $\mathcal{X}$ is bounded, there exists a $B>0$ depending on the norm used on $\mathcal{X}.$ In any case $B=d$ is sufficient. To apply Theorem 5.3 we need the Hölder constant of the network to be greater or equal than $2B.$
Appendix A Calculations for Example 3.3

### v2

@@ A.3 Proofs for Section 4
$=\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\frac{1}{L}\mathsf{W}_{L}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\frac{1}{L}\mathsf{W}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{\hat{G}_{n}(Z)})+\frac{2}{L}\inf_{W\in\mathcal{W}}\sup_{W^{\prime}\in\operatorname{Lip}(L,2B)}\|W-W^{\prime}\|_{\infty}$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\frac{1}{L}\mathsf{W}_{\mathcal{W}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})+\frac{2}{L}\inf_{W\in\mathcal{W}}\sup_{W^{\prime}\in\operatorname{Lip}(L,2B)}\|W-W^{\prime}\|_{\infty}$
$\leq\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})+\frac{1}{L}\mathsf{W}_{\mathcal{H}^{\alpha}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})+\frac{2}{L}\inf_{W\in\mathcal{W}}\sup_{W^{\prime}\in\operatorname{Lip}(L,2B)}\|W-W^{\prime}\|_{\infty}$
The bound on $\mathsf{W}_{\mathcal{H}^{\alpha}}(\mathbb{P}_{n},\mathbb{P}^{G(Z)})$ depending on the intrinsic dimension $d^{*}$ was already derived in Theorem 5.2 (starting with Equation (15)). The bound on $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}_{n})$ depending on the intrinsic dimension $d^{*}$ was already derived in Corollary 4.2.
A.7 Calculations for Example 3.3